OpenStudy (anonymous):

express the given quantity as a single logarithm (1/2)ln(x) - 5ln(x^9 +5)

rishavraj (rishavraj):

(m)ln(x) = ln(x)^m ln x - ln y = ln(x/y)

OpenStudy (anonymous):

ln(x)^1/2 - ln(x^9+5)^5= ln((x)^1/2)/(x^9+5)^5)

OpenStudy (anonymous):

is that right?

rishavraj (rishavraj):

yup..:))))))

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

thanks a lot!

rishavraj (rishavraj):

yeah...nd u r most welcome @swift_13